# Dynamic Matroids: Base Packing and Covering

### v1

@@ 1.2 Organization
In Section˜2, we show how to maintain a minimum weight base dynamically. In Section˜3, we show that dynamic base collections can be used to obtain the dynamic packing and covering results. Hereto, we show the structural results on base packing, base covering, and their connection to base collections in Appendix˜A). And finally, we show that greedy base collections approximate the ideal base collections and how to maintain them in Appendix˜B.
@@ 2 Maintaining a Dynamic Minimum Weight Base
Given a universe $E$, and weight function $w\colon E\to[1,\dots,W]$, a minimum weight base is a base $B$ of minimum total weight $w(B)=\sum_{e\in B}w(e)$. Note that if the weights are unique, the minimum weight base is unique. We can force this by lexicographical tie-breaking, essentially changing the maximum weight to $nW$. The minimum weight base of a matroid can be computed using a simple greedy approach as shown in [gale68, Edmonds71].
For a dynamic minimum weight base, the goal is to maintain a minimum weight base under element insertions and deletions. More generally, we can also allow for weight increases or decreases, but this can also be modeled by deleting the element and inserting it again with the new weight.
@@
We introduce the following notation: $B_{\leq t}:=\{f\in B:w(f)\leq t\}$ and similarly $E_{\leq t}:=\{f\in E:w(f)\leq t\}$}. Let us first consider element insertion. Suppose $e$ is inserted, finding the circuit in $B+e$ can be costly. However, we just need to find the element $f$ of highest weight on this circuit (possibly $e$ itself if it was inserted with a high weight). Formally, we do this as follows.
@@
Note that now $f$ is the element of highest weight on the circuit of $B+e$. Next, we show correctness. We assume that before the insertion $B$ was a base with minimum weight. Note that all subsets of $E$ that were previously independent, are still independent after the addition of $e$ to the matroid. The algorithm outputs the element $f$. In case (a) $w(f)>w(e)$ the maintained base $B^{\prime}$ is set to Bf+eB-f+e, otherwise in case (b) $w(f)<w(e)$ $B^{\prime}=B$ remains unchanged. Now assume towards a contradiction that after the insertion there is another base $B^{\prime\prime}$ such that $w(B^{\prime\prime})<w(B^{\prime})$. Note that $e\in B^{\prime\prime}$, otherwise $B$ would not have been a minimum weight base. Further, there is an element $e^{\prime}$ in C(B+e)eC(B+e)-e with $B^{\prime\prime}-e+e^{\prime}\in\mathcal{I}$. To see this, let $E_{c}$ be the set of elements of $c-B^{\prime\prime}$ for a circuit $c$. If adding an element of $E_{C(B+e)}$ to $B^{\prime\prime}$ results in a circuit containing $e$, we are done. Otherwise, we know that $e\not\in C(B^{\prime\prime}+e^{\prime\prime})$ for any $e^{\prime\prime}\in E_{C(B+e)}$. Considering such an element $e_{1}$ from $E_{C(B+e)}$, by Lemma˜2.1 we know that there is a circuit $c_{1}\subseteq(C(B^{\prime\prime}+e_{1})+C(B+e))-e_{1}$ with $e\in c_{1}$. If $|E_{c_{1}}|=1$ we are done, otherwise (since $c_{1}\subseteq B^{\prime\prime}+E_{C(B+e)}-e_{1}$) we can continue by removing the next element $e_{2}\in c_{1}\setminus B^{\prime\prime}$ in the same way to find a circuit $c_{2}\subseteq(C(B^{\prime\prime}+e_{2})+c_{1})-e_{2}\subseteq B^{\prime\prime}+E_{c_{1}}-e_{2}$ with $e\in c_{2}$. We can continue in the same way, further restricting the number of elements in the next circuit that are not from $B^{\prime\prime}$ until we get a circuit $c^{\prime}\subseteq B^{\prime\prime}+e_{j}$ with $e\in c^{\prime}$. Hence, for the element $e_{j}\in C(B+e)-e$ we have $B^{\prime\prime}-e+e_{j}$ is a base even before the insertion of $e$. Since $e_{j}\in C(B+e)$, we have $w(e_{j})\leq w(f)$. Now, in case (a) where the maintained base was set to Bf+eB-f+e we get
@@
Next, we consider the deletion of an element $e$, which restricts the matroid to EeE-e. The update is similar to before.
@@
For the number of queries, note that we can find the minimum with a binary search over the space of all weight $1,\dots,Wn]$, where the factor $n$ compared to maximum weight $W$ comes from the fact that we add the lexicographical order. ∎
We use Proposition˜1.1 to maintain a greedy base collection. The details are in Appendix˜B.
@@ 3 Dynamic Packing and Covering
In this section, we combine the structural results from Appendix˜A with the results on (dynamic) base collections from Sections˜2 and B to obtain efficient algorithms for base packing and base covering. Some of the theorems follow directly by maintaining a greedy base collection of a large enough size. For other results, we use additional techniques specific to the dynamic challenges at hand.
@@ 3.1 Dynamic Matroid Packing
We maintain a base collection $\mathcal{B}$ of $\Theta\left(\frac{\Phi_{\max}\log(n)}{\varepsilon^{2}}\right)$ greedy bases and maintain the maximum $\ell^{\mathcal{B}}(e)$ using a max-heap directly giving the approximation.
@@
From Lemma˜1.6 we know that if we maintain a base collection of size $\Omega(\Phi\log n/\varepsilon^{\prime 2})$, $\ell^{\mathcal{B}}(e)\geq\ell^{*}(e)-\frac{\varepsilon^{\prime}}{\Phi}$ for all $e\in E$. This gives
@@
Next, we consider the number of queries per update. As updating $\mathcal{B}$ can require $|\mathcal{B}|^{2}$ updates to a greedy base in $\mathcal{B}$ maintaining the base collection takes $O\left(\frac{\Phi_{\max}^{2}\log^{2}(n)\log(nW)}{\varepsilon^{4}}\right)$ queries for any update (worst-case), see Lemma˜1.7. Since $W\leq|\mathcal{B}|=O(\operatorname{poly}n)$, we can write $\log(Wn)=O(\log n)$ and hence we have $O\left(\frac{\Phi_{\max}^{2}\log^{3}(n)}{\varepsilon^{4}}\right)$ queries for any update (worst-case).
@@
bases. We want to exploit this fact, together with the fact that we only need to consider the first $\Theta(\Phi\log n/\varepsilon^{2})$ bases of $\mathcal{B}$ when the base packing number is $\Phi$.
To exploit this, we partition the bases of $\mathcal{B}$ into buckets of size $2^{i}$. At any point in time when the packing number is $\Phi$ and $3\Phi\log n/\varepsilon^{2}\in(2^{i},2^{i+1}]$, we can process an update $e$ in the bases of the buckets efficiently. Denote $\Phi$ for the packing number before the update and $\Phi^{\prime}$ for the packing number after the update.
First, we consider an insertion such that $\Phi^{\prime}\geq\Phi$. We have $\Phi^{\prime}\leq\Phi+1\leq 2\Phi$. Note that after insertion, $e$ will be contained in at most $2\cdot 2^{i+2}/\Phi^{\prime}=\Theta(\log n/\varepsilon^{2})$ bases. When $e$ is part of some base $B_{j}$, this can lead to at most $|\mathcal{B}|$ recourse in all of $\mathcal{B}$. So inserting $\mathcal{B}$ takes at most $|\mathcal{B}|+\Theta(|\mathcal{B}|\log n/\varepsilon^{2})=\Theta(|\mathcal{B}|\log n/\varepsilon^{2})$, where the first factor $\mathcal{B}$ is due to determining which bases to insert $e$ in.
@@
We are left with two cases: the update is an insertion and $\Phi^{\prime}<\Phi$, or the update is a deletion and $\Phi^{\prime}>\Phi$. Note that in graphs, this does not appear: such updates change the graph from disconnected to connected and the packing number of a disconnected graph is 0.
@@ 3.2.1 Deterministic
Analogously to the packing case, we maintain a base collection $\mathcal{B}$ of $\Theta\left(\frac{\beta_{\max}\log(n)}{\varepsilon^{2}}\right)$ greedy bases and maintain the minimum $\ell^{\mathcal{B}}(e)$ using a heap directly giving the approximation.
@@
Maintaining the base collection takes $O\left(\frac{\beta_{\max}^{2}\log^{3}(n)}{\varepsilon^{4}}\right)$ queries for any update (worst-case), see Lemma˜1.7. Since $W\leq|\mathcal{B}|=O(\operatorname{poly}n)$, we can write $\log(Wn)=O(\log n)$ and hence we have $O\left(\frac{\beta_{\max}^{2}\log^{3}(n)}{\varepsilon^{4}}\right)$ queries for any update (worst-case). ∎
@@ 3.2.2 Oblivious adversary
We create $O(\log n)$ matroids $\mathcal{M}_{i}$ by sampling each element with probability $p_{i}=\frac{24c\log n}{2^{i}\varepsilon^{2}}$} (for some constant $c\geq 2$ to be set later and $i$ such that $p_{i}<1$) to get the respective universe. $\mathcal{M}_{i}$ is then given by the restriction of the original matroid to the sampled elements $\mathcal{M}_{i}=\mathcal{M}|E_{i}$. Now we want to show that if $\beta\in[2^{i-1},2^{i+2})$ then the $\frac{1}{p_{i}}\beta_{i}$ is a $(1\pm\varepsilon)$-approximation of $\beta$. We show this in two parts.
@@
Hence, we get that $\beta_{i}\leq(1+\varepsilon)\beta p_{i}\leq 2^{3}\frac{24c\log n}{\varepsilon^{2}}=O\left(\frac{\log n}{\varepsilon^{2}}\right)$ when $\beta\leq 2^{i+2}$. The algorithm maintains a greedy base collection of size $\Theta\left(\frac{\log^{2}n}{\varepsilon^{4}}\right)$ for each of the $O(\log n)$ matroids $\mathcal{M}_{i}$ and one for $G$ itself. Whenever $\beta\in[2^{i-1},2^{i+2})$ the collection corresponding to $\mathcal{M}_{i}$ gives the correct approximation. All other collections might give values that differ drastically from $\beta$ and are simply disregarded. At any point in time to know which $\mathcal{M}_{i}$ gives the correct value it suffices to consider the estimate from the previous update. This is because if the old estimate $\frac{\beta_{i}}{p_{i}}$ was in $2^{i},2^{i+1})$, then after the update we have that $\beta<(1+\varepsilon)\frac{\beta_{i}}{p_{i}}+1\leq 2^{i+2}$ and similarly $\beta\geq(1-\varepsilon)\frac{\beta_{i}}{p_{i}}-1\geq 2^{i-1}$. If the current estimate is in $\Theta(\log n/\varepsilon^{2})$ the estimate on the original matroid is considered, otherwise the interval $2^{i},2^{i+1})$ that contains the previous estimate determines the base collection $\mathcal{M}_{i}$ that is used for the estimate after the next update. Now it remains to analyze the runtime. For each matroid $\mathcal{M}_{i}$ we maintain a greedy base collection of size $O\left(\frac{\log^{2}n}{\varepsilon^{4}}\right)$.
@@
Appendix A Packing, Covering, and Base Collections
@@ A.1 Base Packing
We prove the two inequalities $\leq"$ and $\geq"$, starting with the former, which follows directly without using Theorem˜A.1.
$\leq"$. Let $A^{*}\subseteq E$ be such that $\frac{|A^{*}|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})}=\min\limits_{\begin{subarray}{c}A\subseteq E\ \rm{s.t.}\\ \operatorname{rk}(\overline{A})<\operatorname{rk}(E)\end{subarray}}\frac{|A|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{A})}$. This means that any base $B$ needs to contain $\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})$ elements from $A^{*}$. So for any base collection $\mathcal{B}$, we have $\sum_{e\in A^{*}}\ell^{\mathcal{B}}(e)\geq\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})$. This means that on average for $e\in A^{*}$ we have $\ell^{\mathcal{B}}(e)\geq\frac{\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})}{|A^{*}|}$. In particular $\max_{e\in E}\ell^{\mathcal{B}}(e)\geq\max_{e\in A^{*}}\ell^{\mathcal{B}}(e)\geq\frac{\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})}{|A^{*}|}$. Equivalently, we get that for any base collection $\mathcal{B}$
@@
Now suppose that $\min\limits_{\begin{subarray}{c}A\subseteq E\ \rm{s.t.}\\ \operatorname{rk}(\overline{A})<\operatorname{rk}(E)\end{subarray}}\frac{|A|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{A})}=k$, by Theorem˜A.1 this means that we can pack $k$ disjoint bases. Let $\mathcal{B}$ be the base collection consisting of these $k$ bases. Then for each element $e\in B\in\mathcal{B}$, we have $\ell^{\mathcal{B}}(e)=1/k$. For other elements, we have $\ell^{\mathcal{B}}(e)=0$. Hence we have $\max_{e\in E}\ell^{\mathcal{B}}(e)=1/k$. Equivalently
@@
Consider an arbitrary step $i$ where the matroid $\mathcal{M}_{i}=(E,\mathcal{I})$ is processed and let in the following $\Phi:=\Phi_{\mathcal{M}_{i}}$. Let further $A_{0}$ be a set that gives $|A_{0}|/(\operatorname{rk}(E)-\operatorname{rk}(E\setminus A_{0}))$. Assume that there is a set $A_{1}\subset E\setminus A_{0}$ such that $\frac{|A_{1}|}{\operatorname{rk}(E\setminus A_{0})-\operatorname{rk}((E\setminus A_{0})\setminus A_{1})}<\Phi$, or equivalently $\operatorname{rk}(E\setminus A_{0})-\operatorname{rk}((E\setminus A_{0})\setminus A_{1})>|A_{1}|/\Phi$. By definition of $\Phi$, we also have $\frac{|A_{0}|}{\operatorname{rk}(E)-\operatorname{rk}(E\setminus A_{0})}=\Phi$, or equivalently: $\operatorname{rk}(E)-\operatorname{rk}(E\setminus A_{0})=|A_{0}|/\Phi$.
@@
Suppose we have two sets $S$ and $T$ such that
@@
By Lemma˜A.2, we already know that it is at least $\Phi$. So it remains to show that it is at most $\Phi$. For contradiction, assume

### v2

@@ 1.2 Organization
In Section˜2, we show how to maintain a minimum weight base dynamically. In Section˜3, we show that dynamic base collections can be used to obtain the dynamic packing and covering results. Hereto, we show the structural results on base packing, base covering, and their connection to base collections in Section˜4. And finally, we show that greedy base collections approximate the ideal base collections and how to maintain them in Section˜5.
@@ 2 Maintaining a Dynamic Minimum Weight Base
Given a universe $E$, and weight function $w\colon E\to[1,\dots,W]$, a minimum weight base is a base $B$ of minimum total weight $w(B)=\sum_{e\in B}w(e)$. Note that if the weights are unique, the minimum weight base is unique. For simplicity, we would like the weights to be unique without increasing the maximum weight. The minimum weight base of a matroid can be computed using a simple greedy approach as shown in [gale68, Edmonds71]. Here, the exact weights of the elements are never taken into account. It suffices to consider the order on the elements implied by their weights. Thus we can essentially re-weight the elements at each update, assigning each element a unique weight in $1,\dots,n]$ respecting, firstly, the weights given by $w$, and secondly, in case of equal weight, the lexicographic order of the unique ids. [^6] Note that after each update, many weights can change. However, this does not require additional queries, so does not affect our query complexity. For a dynamic minimum weight base, the goal is to maintain a minimum weight base under element insertions and deletions. More generally, we can also allow for weight increases or decreases, but this can also be modeled by deleting the element and inserting it again with the new weight.
@@
We introduce the following notation: the sets $B_{\leq t}:=\{f\in B:w(f)\leq t\}$ and similarly $E_{\leq t}:=\{f\in E:w(f)\leq t\}$}. Let us first consider element insertion. If the insertion increases the rank of the matroid, we can simply add the new element to $B$ to get a minimum weight base for the new matroid. Otherwise, suppose $e$ is inserted, finding the circuit in $B+e$ can be costly. However, we just need to find the element $f$ of highest weight on this circuit (possibly $e$ itself if it was inserted with a high weight). Formally, we do this as follows.
@@
Note that now $f$ is the element of highest weight on the circuit of $B+e$. Next, we show that we can use this method to maintain a minimum base correctly over a sequence of insertions that do not increase the rank of the matroid. We assume that before the insertion $B$ was a base with minimum weight. Note that all subsets of $E$ that were previously independent, are still independent after the addition of $e$ to the matroid. The algorithm outputs the element $f$. In case (a) $w(f)>w(e)$ the maintained base $B^{\prime}$ is set to Bf+eB-f+e, otherwise in case (b) $w(f)<w(e)$ $B^{\prime}=B$ remains unchanged. Now assume towards a contradiction that after the insertion there is another base $B^{\prime\prime}$ such that $w(B^{\prime\prime})<w(B^{\prime})$. Note that $e\in B^{\prime\prime}$, otherwise $B$ would not have been a minimum weight base. Further, there is an element $e^{\prime}$ in C(B+e)eC(B+e)-e with $B^{\prime\prime}-e+e^{\prime}\in\mathcal{I}$. To see this, let $E_{c}$ be the set of elements of $c-B^{\prime\prime}$ for a circuit $c$. If adding an element of $E_{C(B+e)}$ to $B^{\prime\prime}$ results in a circuit containing $e$, we are done. Otherwise, we know that $e\not\in C(B^{\prime\prime}+e^{\prime\prime})$ for any $e^{\prime\prime}\in E_{C(B+e)}$. Considering such an element $e_{1}$ from $E_{C(B+e)}$, by Lemma˜2.1 we know that there is a circuit $c_{1}\subseteq(C(B^{\prime\prime}+e_{1})+C(B+e))-e_{1}$ with $e\in c_{1}$. If $|E_{c_{1}}|=1$ we are done, otherwise (since $c_{1}\subseteq B^{\prime\prime}+E_{C(B+e)}-e_{1}$) we can continue by removing the next element $e_{2}\in c_{1}\setminus B^{\prime\prime}$ in the same way to find a circuit $c_{2}\subseteq(C(B^{\prime\prime}+e_{2})+c_{1})-e_{2}\subseteq B^{\prime\prime}+E_{c_{1}}-e_{2}$ with $e\in c_{2}$. We can continue in the same way, further restricting the number of elements in the next circuit that are not from $B^{\prime\prime}$ until we get a circuit $c^{\prime}\subseteq B^{\prime\prime}+e_{j}$ with $e\in c^{\prime}$. Hence, for the element $e_{j}\in C(B+e)-e$ we have $B^{\prime\prime}-e+e_{j}$ is a base even before the insertion of $e$. Since $e_{j}\in C(B+e)$, we have $w(e_{j})\leq w(f)$. Now, in case (a) where the maintained base was set to Bf+eB-f+e we get
@@
Next, we consider the deletion of an element $e$, which restricts the matroid to EeE-e. If the deletion decreases the rank of the matroid, the element cannot be replaced and $B\setminus\{e\}$ gives the new minimum weight base. Otherwise, we need to find an element $f$ of minimum weight such that $B\cup\{f\}$ contains a circuit in the matroid before the deletion. Similar to an insertion, the approach is as follows.
@@
For the number of queries, note that we can find the minimum with a binary search over the space of all adjusted weights $1,\dots,n]$. ∎
We use Proposition˜1.1 to maintain a greedy base collection. The details are in Section˜5.
@@ 3 Dynamic Packing and Covering
In this section, we combine the structural results from Section˜4 with the results on (dynamic) base collections from Sections˜2 and 5 to obtain efficient algorithms for base packing and base covering. Some of the theorems follow directly by maintaining a greedy base collection of a large enough size. For other results, we use additional techniques specific to the dynamic challenges at hand.
@@ 3.1 Dynamic Matroid Packing
We maintain a base collection $\mathcal{B}$ of $\Theta\left(\frac{\Phi_{\max}\log n}{\varepsilon^{2}}\right)$ greedy bases and maintain the maximum $\ell^{\mathcal{B}}(e)$ using a max-heap directly giving the approximation.
@@
From Lemma˜1.6 we know that if we maintain a base collection of size $\Omega(\Phi\log n/\varepsilon^{\prime 2})$, $\ell^{\mathcal{B}}(e)\geq\ell^{*}(e)-\frac{\varepsilon^{\prime}}{\Phi}$ for all $e\in E$ for an $\varepsilon^{\prime}>0$ to be defined later, this gives
@@
Next, we consider the number of queries per update. As updating $\mathcal{B}$ can require $|\mathcal{B}|^{2}$ updates to a greedy base in $\mathcal{B}$ maintaining the base collection takes $O\left(\frac{\Phi_{\max}^{2}\log^{3}(n)}{\varepsilon^{4}}\right)$ queries for any update (worst-case), see Lemma˜1.7.
@@
bases. We want to exploit this fact, together with the fact that we only need to consider the first $\Theta(\Phi\log n/\varepsilon^{2})$ bases of $\mathcal{B}$ when the base packing number is $\Phi$ by Lemma˜1.6. To exploit this, we partition $\mathcal{B}$ into buckets, where the $i$th bucket contains $2^{i}$ bases, with $i=0,\dots,\Theta(\log|\mathcal{B}|)$. Intuitively, at any point in time when the packing number is $\Phi$ and $3\Phi\log n/\varepsilon^{2}\in(2^{i},2^{i+1}]$, we can process an update $e$ in the bases of the first $i+1$ buckets efficiently – which is we all we need to output our estimate of $\Phi$. More formally, denote $\Phi$ for the packing number before the update and $\Phi^{\prime}$ for the packing number after the update.
First, we consider an insertion such that $\Phi^{\prime}\geq\Phi$. We have $\Phi^{\prime}\leq\Phi+1\leq 2\Phi$. Note that after insertion, $e$ will be contained in at most $2\cdot 2^{i+2}/\Phi^{\prime}=\Theta(\log n/\varepsilon^{2})$ bases. When $e$ is part of some base $B_{j}$, this can lead to at most $|\mathcal{B}|$ recourse in all of $\mathcal{B}$ (see Lemma˜1.7 for the proof). So inserting $\mathcal{B}$ takes at most $|\mathcal{B}|+\Theta(|\mathcal{B}|\log n/\varepsilon^{2})=\Theta(|\mathcal{B}|\log n/\varepsilon^{2})$, where the first factor $\mathcal{B}$ is due to determining which bases to insert $e$ in.
@@
We are left with two cases: the update is an insertion and $\Phi^{\prime}<\Phi$, or the update is a deletion and $\Phi^{\prime}>\Phi$. Note that for tree-packing in graphs, this does not appear: such updates change the graph from disconnected to connected and the packing number of a disconnected graph is 0.
@@ 3.2.1 Deterministic
Analogously to the packing case, we maintain a base collection $\mathcal{B}$ of $\Theta\left(\frac{\beta_{\max}\log n}{\varepsilon^{2}}\right)$ greedy bases and maintain the minimum $\ell^{\mathcal{B}}(e)$ using a heap directly giving the approximation.
@@
Maintaining the base collection takes $O\left(\frac{\beta_{\max}^{2}\log^{3}(n)}{\varepsilon^{4}}\right)$ queries for any update (worst-case), see Lemma˜1.7. ∎
@@ 3.2.2 Oblivious adversary
We create $O(\log n)$ matroids $\mathcal{M}_{i}$ by sampling each element with probability $p_{i}=\frac{24c\log n}{2^{i}\varepsilon^{2}}$} (for some constant $c\geq 2$ to be set later and $i$ such that $p_{i}<1$) to get the respective universe $E_{i}$. $\mathcal{M}_{i}$ is then given by the restriction of the original matroid to the sampled elements $\mathcal{M}_{i}=\mathcal{M}|E_{i}$. Now we want to show that if $\beta\in[2^{i-1},2^{i+2})$ then $\frac{1}{p_{i}}\beta_{i}$ is a $(1\pm\varepsilon)$-approximation of $\beta$. We show this in two parts.
@@
Hence, we get that $\beta_{i}\leq(1+\varepsilon)\beta p_{i}\leq 2^{3}\frac{24c\log n}{\varepsilon^{2}}=O\left(\frac{\log n}{\varepsilon^{2}}\right)$ when $\beta\leq 2^{i+2}$. The algorithm maintains a greedy base collection of size $\Theta\left(\frac{\log^{2}n}{\varepsilon^{4}}\right)$ for each of the $O(\log n)$ matroids $\mathcal{M}_{i}$ and one for $\mathcal{M}$ itself. Whenever $\beta\in[2^{i-1},2^{i+2})$ the collection corresponding to $\mathcal{M}_{i}$ gives the correct approximation. All other collections might give values that differ drastically from $\beta$ and are simply disregarded. At any point in time to know which $\mathcal{M}_{i}$ gives the correct value it suffices to consider the estimate from the previous update. This is because if the old estimate $\frac{\beta_{i}}{p_{i}}$ was in $2^{i},2^{i+1})$, then after the update we have that $\beta<(1+\varepsilon)\frac{\beta_{i}}{p_{i}}+1\leq 2^{i+2}$ and similarly $\beta\geq(1-\varepsilon)\frac{\beta_{i}}{p_{i}}-1\geq 2^{i-1}$. If the current estimate is in $\Theta(\log n/\varepsilon^{2})$ the estimate on the original matroid is considered, otherwise the interval $2^{i},2^{i+1})$ that contains the previous estimate determines the base collection $\mathcal{M}_{i}$ that is used for the estimate after the next update. Now it remains to analyze the runtime. For each matroid $\mathcal{M}_{i}$ we maintain a greedy base collection of size $O\left(\frac{\log^{2}n}{\varepsilon^{4}}\right)$.
@@
4 Packing, Covering, and Base Collections
@@ 4.1 Base Packing
We prove the two inequalities $\leq"$ and $\geq"$, starting with the former, which follows directly without using Theorem˜4.1.
$\leq"$. Let $A^{*}\subseteq E$ be such that $\frac{|A^{*}|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})}=\min\limits_{\begin{subarray}{c}A\subseteq E\ \rm{s.t.}\\ \operatorname{rk}(\overline{A})<\operatorname{rk}(E)\end{subarray}}\frac{|A|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{A})}$. Now we note that any base $B$ needs to contain $\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})$ elements from $A^{*}$. So for any base collection $\mathcal{B}$, we have $\sum_{e\in A^{*}}\ell^{\mathcal{B}}(e)\geq\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})$. This means that on average for $e\in A^{*}$ we have $\ell^{\mathcal{B}}(e)\geq\frac{\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})}{|A^{*}|}$. In particular $\max_{e\in E}\ell^{\mathcal{B}}(e)\geq\max_{e\in A^{*}}\ell^{\mathcal{B}}(e)\geq\frac{\operatorname{rk}(E)-\operatorname{rk}(\overline{A^{*}})}{|A^{*}|}$. Equivalently, we get that for any base collection $\mathcal{B}$
@@
Now suppose that $\min\limits_{\begin{subarray}{c}A\subseteq E\ \rm{s.t.}\\ \operatorname{rk}(\overline{A})<\operatorname{rk}(E)\end{subarray}}\frac{|A|}{\operatorname{rk}(E)-\operatorname{rk}(\overline{A})}=k$, by Theorem˜4.1 this means that we can pack $k$ disjoint bases. Let $\mathcal{B}$ be the base collection consisting of these $k$ bases. Then for each element $e\in B\in\mathcal{B}$, we have $\ell^{\mathcal{B}}(e)=1/k$. For other elements, we have $\ell^{\mathcal{B}}(e)=0$. Hence we have $\max_{e\in E}\ell^{\mathcal{B}}(e)=1/k$. Equivalently
@@
Consider an arbitrary step $i$ where the matroid $\mathcal{M}_{i}=(E,\mathcal{I})$ is processed and let in the following $\Phi:=\Phi_{\mathcal{M}_{i}}$. Let further $A_{0}$ be a set that gives $\Phi=|A_{0}|/(\operatorname{rk}(E)-\operatorname{rk}(E\setminus A_{0}))$. Assume that there is a set $A_{1}\subset E\setminus A_{0}$ such that $\frac{|A_{1}|}{\operatorname{rk}(E\setminus A_{0})-\operatorname{rk}((E\setminus A_{0})\setminus A_{1})}<\Phi$, or equivalently $\operatorname{rk}(E\setminus A_{0})-\operatorname{rk}((E\setminus A_{0})\setminus A_{1})>|A_{1}|/\Phi$. By definition of $\Phi$, we also have $\frac{|A_{0}|}{\operatorname{rk}(E)-\operatorname{rk}(E\setminus A_{0})}=\Phi$, or equivalently: $\operatorname{rk}(E)-\operatorname{rk}(E\setminus A_{0})=|A_{0}|/\Phi$.
@@
Suppose we have two sets $S$ and $T$ such that $T\not\subseteq S$ and
@@
First, we note that $\operatorname{rk}(E\setminus S)-\operatorname{rk}((E\setminus S)\setminus(T\cap(E\setminus S)))=\operatorname{rk}(E\setminus S)-\operatorname{rk}((E\setminus S)\cap(E\setminus T))\neq 0$ as otherwise we would get that $\operatorname{rk}((E\setminus S)\cup(E\setminus T))=\operatorname{rk}(E\setminus(S\cap T))\leq\operatorname{rk}(E\setminus T)$. Then it would follow that $\frac{|T\cap S|}{\operatorname{rk}(E)-\operatorname{rk}(E\setminus(T\cap S))}<\Phi$ which contradicts the definition of $\Phi$. By Lemma˜4.2, we already know that it is at least $\Phi$. So it remains to show that it is at most $\Phi$. For contradiction, assume